# Efficient detection of entanglement by stimulated disentanglement

### v2

@@
Efficient detection of entanglement by stimulated disentanglement
@@
Standard detection of entanglement relies on local measurements of the individual particles, evaluating their correlations in post-processing. For time-energy entangled photons, either times $(t_{1},t_{2})$ or energies $(E_{1},E_{2})$ are measured, but not both due to the mutual quantum uncertainty, providing only partial information of the entanglement. In principle, a global detector could recover the complete information of entanglement in a single shot if it could measure the combined correlated variables $(t_{1}-t_{2})$ and $(E_{1}+E_{2})$ without measuring the individual energies or times. Such a global measurement is possible using the reverse disentangling interaction, like sum-frequency generation (SFG), but nonlinear interactions at the single-photon level are exceedingly inefficient. We overcome this barrier by stimulating the nonlinear SFG interaction with a strong pump, thereby measuring both the energy-sum (SFG spectrum) and the time-difference (response to group delay/dispersion) simultaneously and efficiently. We generate biphotons with extreme time-energy entanglement (octave-spanning spectrum of 113THz) and measure a relative uncertainty between time-difference and energy-sum of $\Delta(t_{1}-t_{2})\Delta(E_{1}+E_{2})\approx\!2\!\cdot\!10^{-13}h$, violating the classical bound by more than 12 orders of magnitude. The presented coherent SFG dramatically enhances the detection SNR compared to standard methods since it ideally rejects erroneous coincidences in both time and energy, paving the way for sensing applications, such as quantum illumination (radar) and more.
@@ 1 Introduction
Entanglement of conjugate variables, such as position and momentum, or energy and time, is a fundamental aspect of quantum mechanics. The profound nature of entanglement lies in the ability of quantum systems to exhibit perfect non-local correlations between the properties of two or more particles, even when the local values of those properties remain fundamentally undetermined and their simultaneous local measurement is excluded by quantum uncertainty. This paradoxical feature, first recognized by Einstein, Podolsky, and Rosen [1], eventually laid the foundation for our modern understanding of quantum physics, highlighting the inherent non-locality that characterizes quantum correlations. Today, entanglement is a crucial resource in quantum computing [2, 3, 4], communication [5, 6, 7], and metrology [8, 9, 10], where harnessing these correlations across different variables enables significant advancements in processing and transmission of quantum information [11, 12].
@@
The generation of entanglement relies on an entangling interaction between the parties, which for photons inherently entails a nonlinearity, such as three- or four-wave mixing. A well-known example is the generation of entangled photons by spontaneous parametric down-conversion (SPDC) in a three-wave mixing medium, where a pump ($p$) photon from an intense laser splits into a pair of lower-energy photons, named signal ($s$) and idler ($i$) [13, 14]. These generated biphotons inherit the properties of the pump, indicating that the sum of all their preserved properties, like energies (frequencies), linear momenta ($k$-vectors), angular momenta (transverse modes) or spins (polarizations) equal those of the original pump, but the properties of each of the photons are not defined and can span a range of possible values. In this paper, we focus on time-energy entangled pairs of photons that are generated by broadband spontaneous parametric down-conversion from a narrowband pump laser: The total energy is fixed ($\nu_{s}+\nu_{i}=\nu_{p}$), as dictated by the pump, but the individual photon energies vary across a very wide bandwidth, limited only by the phase-matching of the nonlinear medium, which can span up to an octave of bandwidth ($\Delta\nu\!\approx\!113$THz in our experiment). Conversely, the nonlinear generation of the pair is “instantaneous”, indicating that the arrival times of the photons are tightly correlated down to the Fourier limit $t_{s}\!-\!t_{i}\!\leq\!1\!/\!\Delta\nu$ ($\approx\!9$ fs in our experiment).
@@
While generation of extreme entanglement is rather simple, its unambiguous characterization is a challenge due to the collision between the non-local nature of entanglement and the constraints on local measurements due to quantum uncertainty. For example, if we measure locally the energies of both photons $E_{s},E_{i}$ to obtain the energy-correlation, uncertainty precludes any information on the time-correlation. However simultaneous global measurement of both $E_{s}\!+\!E_{i}$, $t_{s}\!-\!t_{i}$ is allowed by uncertainty if the individual energies or times are not measured, thereby uncovering the complete correlation information. If such global observation of both time-difference and energy-sum can be realized efficiently, it will provide a compelling proof of (non-classical) entanglement, since classically, the mutual uncertainty of time-difference and energy-sum is bounded by [15]
@@
Since efficient global detection was not available so far, the standard characterization of entanglement relies on correlations of local measurements. In discrete-variable systems (such as polarization-entangled photons), Bell’s inequalities utilize measurements in carefully selected bases to quantify the quantum violation of the classical correlation bounds [16, 17]. However, the direct extension of Bell’s inequalities to continuous-variable (CV) systems remains an ongoing challenge, both theoretically and experimentally. Many important techniques were proposed, including continuous variables (CV) Bell inequalities, e.g. [18, 19, 20, 21, 22], phase-space reconstructions [23, 24], variance-based inequalities (i.e. the Simon-Duan criterion and its generalization to higher moments [25, 26, 27, 28, 29], and more general Gaussian-state criteria [30], but no single method can universally and conclusively detect all forms of (CV) entanglement. Moreover, some of the above methods are challenging in practice, involving complex preparation of sophisticated measurements, and may result in very small Bell violations [31]. Moreover, limitations introduced by detector inefficiencies, optical losses, or finite sampling rates can obscure genuine quantum correlations. Nevertheless, various techniques were developed to witness and quantify time-energy entanglement, as summarized in Fig. 1a, including coincidence detection in time with Hong-Ou-Mandel (HOM) interference [32], spectral-correlation measurement using SU(1,1) interferometry [33, 34] and direct sum-frequency generation (SFG) [35, 36, 37]. While these methods have advanced the field, none of them provided the complete correlation efficiently from both time and energy.
@@
Furthermore, the enhanced SNR of SFG detection directly indicates the time-energy entanglement: While the biphoton amplitude of each individual signal-idler frequency-pair in the spectrum is well below the detection threshold of $1\sqrt{\text{photons}}/\text{mode}$, the coherent sum of all modes across the spectrum can generate a well-detectable SFG signal with practically no noise background due to the ideal rejection of accidental coincidences (as we discuss hereon). Consequently, SFG is recognized as ideal for quantum sensing applications that rely on entanglement, such as quantum radar and imaging with quantum illumination [39, 37]. Unfortunately, the efficiency of direct SFG at the single-photon level is dramatically low (typically $10^{-8}$ or lower). Thus, although SFG of entangled photons was demonstrated to expose the non-classical features of time-energy entangled photons [36]), it cannot form a practical detector for real-world applications. The table in Fig. 1b summarizes the advantages and limitations of the existing methods of entanglement detection.
@@
To overcome the inefficiency barrier, we propose to stimulate the nonlinear interaction in the detection crystal with a strong coherent pump field, which dramatically enhances the SFG efficiency of photon-pairs to near unity [40]. Coherent stimulation—now phase-dependent—underpins nonlinear SU(1,1) interference, where the relative phase between the pump field and biphotons determines the energy flow in the second nonlinear crystal (from biphotons to pump or vice versa). This non-classical SU(1,1) interference is also used for biphoton detection (Fig. 1a) by measuring the spectral intensity of biphotons after the detection crystal. The resulting high-contrast interference fringes reveal the nonlinear pairwise interference of photons, with fringe contrast serving as a non-classical witness [33]. Remarkably, this process remains efficient even in lossy conditions, selectively extracting correlated biphotons. Moreover, SU(1,1) coherence enabled parametric homodyne detection [41], which uses the pump as a common phase reference to measure a specified two-mode quadrature of all signal-idler frequency-pairs across a broad spectrum. And yet, SU(1,1) detection does not capture the complete quantum correlation, since it relies on local measurements of the down-converted photons.
@@
To obtain full correlation information, we return to SFG detection with a twist: coherent SFG (Fig. 1a). This method leverages the near unit-efficiency of SU(1,1) interference while shifting detection from the biphotons back to the pump—measuring the pump’s depletion, gain, or phase shift from the coherent SFG conversion. In doing so, it combines SFG’s global, noise-rejecting nature with the efficiency of SU(1,1) interference.
@@
Experimentally, coherent SFG presents a major challenge, as it aims to measure a small SFG addition / depletion on top of an intense pump beam (of $\sim\!400$mW power), which saturates the detector by orders of magnitude. To overcome this we mix the output beam of pump+SFG with a local oscillator at the pump frequency, whose phase is tuned to destructively interfere with the pump, thereby coherently removing most of the strong pump background from the desired SFG result before the detection. Using coherent seeding of the SFG process and lock-in detection we are able to stabilize the phase of the small SFG signal.
@@
In addition, we present an analytical quantum model (see the Methods section) of coherent SFG that we developed to fit the experimental data. The model provides a closed expression for the SFG amplitude beyond the “undepleted pump” approximation [14] by treating the three-wave mixing process as a perturbative expansion in orders of the spontaneous parametric down conversion (SPDC) amplitude. Our model is applicable in a wide range of parametric gain, from the low gain regime of single entangled photon pairs ($g\!\ll\!1$) up to the multi-photon regime of high quadrature squeezing ($g\!\geq\!1$).
@@ 2 Experiment and Results
The main results are shown in Fig. 3, presenting the correlation in both energy-sum and time-difference (or group delay). The top panel (a) reflects the SFG spectrum of recombination, highlighting the precise match of the sum-energy of the biphotons with the original pump energy and confirming that the process is dominated by the recombination of correlated photons back to the pump frequency $\nu_{SFG}\!=\!\nu_{p}$, while the probability of accidental SFG to other SFG frequencies from non-twin photons vanishes (as theoretically predicted, see below). Specifically, panel (a) shows the RF spectrum of the of the detected pump+SFG intensity, demonstrating a single peak of high contrast ($>\!37$dB) at the lock-in dither frequency $\omega$. The measured intensity modulation is due to the beat of interference between the carrier pump at $\nu_{P}$ and the SFG sidebands at $\nu_{SFG}\!=\!\nu_{P}\!\pm\!\omega$ due to the phase modulation of the SPDC, thereby capturing the frequency-difference between the SFG signal and the pump at Hz-level resolution ($20$Hz width), limited only by the integration time of the RF analyzer. The background of the SFG peak is purely the electronic noise floor of our digital spectrum analyzer (Moku:Lab). The absence of an optical contribution to the background at other frequencies $\nu\!\neq\!\omega$ indicates that the recombination of correlated biphotons dominates and that accidental SFG events are insignificant.
@@
The central panel (b) of Fig. 3 reflects the correlation in time-difference between the photons by observing the dependence of the SFG spectral peak (at the dither frequency $\nu$) on the relative group delay between the signal and the idler photons before entering the SFG crystal. We continuously tune the group delay (and GDD) by varying the amount of material that the beam of biphotons traverses in the spectral-shaper (by shifting the Sapphire prisms in/out). The additional material dispersion imprints a quadratic phase $\phi(\nu)\!=\!\beta_{2}\nu^{2}$ on the broad spectrum of spontaneous biphotons, which in time indicates a group delay (and dispersion) between each signal frequency and its twin idler.; and the SFG signal extracted from the lock-in detection is simply $I_{SFG}\!\sim\!\text{Re}[\alpha_{r}^{*}\sum_{\nu}a_{s}(\nu)a_{i}(-\nu)\mathrm{e}^{\mathrm{i}\phi(\nu)}]$, which reflects a homodyne measurement of the SFG sum over all the signal-idler pairs with $\alpha_{r}$ the coherent amplitude of the pump residue at the Sagnac output (see derivation below). Since even small GDD values alter the phase of the spontaneous SFG field, the variation of the GDD is immediately reflected in the homodyne, causing the sharp peak. The bottom panels (c) and (d) show the spectral SU(1,1) interferograms for two specific GDD values, marked A and B on the central panel (b): A - near GDD compensation (bottom right, blue), where the spectral phase is flat; and B - at a large GDD value (bottom left, violet), where the large quadratic spectral phase is imprinted on the spectral fringes.
@@
Finally, all the GDD-scans of Fig. 4 with various residue phases demonstrate a sharp peak near the zero-GDD point, which is a direct signature of the time correlation of the photon pairs. Specifically, the peak width $\delta\beta\!\approx\!4\pi/\Delta\nu^{2}\approx 30\,\mathrm{fs}^{2}$ is inversely proportional to the biphotons bandwidth of $\Delta\nu\!=\!113$ THz squared (for $\varphi_{r}\!=\!0,\pi$ the peak is split). From a physical point of view, this width is also the GDD tolerance for the photon pairs to keep coherence, which reflects their time correlation of the biphotons at $\Delta t\!=\!1/\Delta\nu\!\approx\!9$ fs. Together with the measured linewidth of $20$Hz for the SFG peak, mentioned above we obtain $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!\approx\!2\!\cdot\!10^{-13}$, showing a giant violation of the classical bound (Eq. 1). The longer-range oscillations in the GDD-scans beyond the central coherent SFG peak are mainly due to the fact that we scan the GDD, rather than the group delay, which indicates that the spectral phase at the center of the spectrum varies quadratically and not linearly across the spectrum.

### v3

# Quantum matched filtering: breaking time-energy separability by 12 orders of magnitude

@@
Quantum matched filtering: breaking time-energy separability by 12 orders of magnitude
@@
Detection of signals buried in noise is the major challenge for sensing. Classically, the optimal detector is a matched filter, whose sensitivity meets the classical limit of correlation between the filter target $T(t_{2})$ and the measured signal within the noise $S(t_{1})$. For classical signals, the correlation is limited by the separability criterion in frequency-time $\Delta(t_{1}\!-\!t_{2})\Delta(\nu_{1}\!+\!\nu_{2})\!\geq\!\frac{1}{2}$. Quantum states, however are not necessarily separable, and the correlation between entangled particles can surpass the classical limits. Specifically, time-energy entangled photons can be simultaneously correlated in time difference $t_{1}-t_{2}$ and frequency sum $\nu_{1}+\nu_{2}$ with no minimum limit of the above inequality, potentially leading to a drastic enhancement of sensitivity for diversified sensing applications. Yet, to enjoy this quantum enhancement, a unique, global detector is needed that can recover the complete information of entanglement in a single shot, i.e. measure the combined correlated variables $(t_{1}-t_{2})$ and $(\nu_{1}+\nu_{2})$ without measuring the individual frequencies or times. Such a global measurement could, in principle, be realized using the reverse disentangling interaction, such as sum-frequency generation (SFG), but nonlinear interactions at the single-photon level have long been prohibitively inefficient, significantly restricting practical implementations. Here we overcome this barrier: We measure simultaneously and efficiently both the frequency-sum (SFG spectrum) and the time-difference (relative group delay/dispersion) by stimulating the SFG recombination with a strong pump. We generate biphotons with extreme time-energy entanglement (octave-spanning spectrum of 113THz) and measure a relative uncertainty between time-difference and frequency-sum of $\Delta(t_{1}\!-\!t_{2})\Delta(\nu_{1}\!+\!\nu_{2})\!\approx\!2\!\cdot\!10^{-13}$, violating the classical separability bound by more than 12 orders of magnitude. Our experiment and supporting theory pave the way for improved sensing applications, such as quantum illumination (radar), and for various continuous-variable communication protocols.
@@ 1 Introduction
Entanglement of conjugate variables, such as position and momentum, or energy and time, is a fundamental aspect of quantum mechanics. The profound nature of entanglement lies in the ability of quantum systems to exhibit perfect nonlocal correlations between the properties of two or more particles, even when the local values of those properties remain fundamentally undetermined and their simultaneous local measurement is excluded by quantum uncertainty. This paradoxical feature, first recognized by Einstein, Podolsky, and Rosen [1], eventually laid the foundation for our modern understanding of quantum physics, highlighting the inherent nonlocality that characterizes quantum correlations. Today, entanglement is a crucial resource in quantum computing [2, 3, 4], communication [5, 6, 7], and metrology [8, 9, 10], where harnessing these correlations across different variables enables significant advancements in processing and transmission of quantum information [11, 12].
@@
The generation of entanglement relies on an entangling interaction between the parties, which for photons inherently entails a nonlinearity, such as three- or four-wave mixing. A well-known example is the generation of entangled photons by spontaneous parametric down-conversion (SPDC) in a three-wave mixing medium, where a pump ($p$) photon from an intense laser splits into a pair of lower-energy photons, named signal ($s$) and idler ($i$) [13, 14]. These generated biphotons inherit the properties of the pump, indicating that the sum of all their preserved properties, like energies (frequencies), linear momenta ($k$-vectors), angular momenta (transverse modes) or spins (polarizations) equal those of the original pump, but the properties of each of the photons are not defined and can span a range of possible values. In this paper, we focus on time-energy entangled pairs of photons that are generated by broadband spontaneous parametric down-conversion from a narrowband pump laser: The frequency (energy) sum is fixed ($\nu_{s}+\nu_{i}=\nu_{p}$), as dictated by the pump, but the individual photon frequencies vary across a very wide bandwidth, limited only by the phase-matching of the nonlinear medium, which can span up to an octave of bandwidth ($\Delta\nu\!\approx\!113$THz in our experiment). Conversely, the nonlinear generation of the pair is “instantaneous”, indicating that the arrival times of the photons are tightly correlated down to the Fourier limit $t_{s}\!-\!t_{i}\!\leq\!1\!/\!\Delta\nu$ ($\approx\!9$ fs in our experiment).
@@
While generation of extreme entanglement is rather simple, its unambiguous characterization is a challenge due to the collision between the nonlocal nature of entanglement and the constraints on local measurements due to quantum uncertainty. For example, if we measure locally the frequencies of both photons $\nu_{s},\nu_{i}$ to obtain the energy-correlation, uncertainty precludes any information on the time-correlation. However, simultaneous global measurement of both $\nu_{s}\!+\!\nu_{i}$, $t_{s}\!-\!t_{i}$ is allowed by uncertainty if the individual frequencies or times are not measured, thereby uncovering the complete correlation information. If such a global observation of both time-difference and frequency-sum can be realized efficiently, it will provide a compelling proof of (non-classical) entanglement, since for classical separable states, the mutual uncertainty of time-difference and energy-sum is bounded by [15]
@@
Since efficient global detection was not available so far, the standard characterization of entanglement relies on correlations of local measurements. In discrete-variable systems (such as polarization-entangled photons), Bell’s inequalities utilize measurements in carefully selected bases to quantify the quantum violation of the classical correlation bounds [16, 17]. However, the direct extension of Bell’s inequalities to continuous-variable (CV) systems remains an ongoing challenge, both theoretically and experimentally. Many important techniques were proposed, including continuous variables (CV) Bell inequalities, e.g. [18, 19, 20, 21, 22], phase-space reconstructions [23, 24], variance-based inequalities (i.e. the Simon-Duan criterion and its generalization to higher moments [25, 26, 27, 28, 29], and more general Gaussian-state criteria [30], but no single method can universally and conclusively detect all forms of (CV) entanglement. Moreover, some of the above methods are challenging in practice, involving complex preparation of sophisticated measurements, and may result in very small Bell violations [31]. In addition, limitations introduced by detector inefficiencies, optical losses, or finite sampling rates can obscure genuine quantum correlations.
Nevertheless, various techniques were developed to witness and quantify time-energy entanglement, as summarized in Fig. 1, including coincidence detection in time with Hong-Ou-Mandel (HOM) interference [32], spectral-correlation measurement using SU(1,1) interferometry [33, 34] and direct sum-frequency generation (SFG) [35, 36, 37] (the table in Fig. 1b summarizes the advantages and limitations of each method). While these methods have advanced the field, none of them provided the complete correlation efficiently from both time and energy.
@@
2 Main Contribution
SFG of broadband PDC forms a powerful optical matched filter for sensing applications: Even classically, the signal and idler fields of broadband PDC are spectral complex conjugates $E_{i}(\nu)=E_{s}^{*}(-\nu)$, which designates the idler (signal) as a matched filter for the signal (idler). This is particularly advantageous for extracting the signal $E_{s}(\nu)$ from a noise environment $n(\nu)$ using broadband SFG against the idler, which implements a spectral correlation between the two [Peer2004CDMA]:
$E_{SFG}(\nu)\propto$ $\!\int\!d\nu^{\prime}\left[E_{s}(\nu^{\prime})\!+\!n(\nu^{\prime})\right]E_{i}(\nu-\nu^{\prime})=$ (2)
$=$ $\!\int\!d\nu^{\prime}E_{s}(\nu^{\prime})E_{s}^{*}(\nu^{\prime}-\nu)+\!\int\!d\nu^{\prime}n(\nu^{\prime})E_{s}^{*}(\nu^{\prime}-\nu)=$ (3)
$=$ $\ S_{corr}(\nu)\!+\!B_{bg}(\nu),$ (4)
where $S_{corr}(\nu=0)=\int d\nu^{\prime}\left|E_{s}(\nu^{\prime})\right|^{2}$ is the coherent correlation peak that appears at the frequency of the original pump and is proportional to the signal’s total power, which is much larger than the incoherent uncorrelated background $B_{bg}(\nu=0)=\int d\nu^{\prime}n(\nu^{\prime})E_{s}^{*}(\nu^{\prime})$ due to the noise. Classically, the noise contribution can therefore be suppressed by the correlation, but cannot be totally eliminated, posing the fundamental limit on the detection sensitivity.
@@ 2 Main Contribution
In contrast, a quantum matched filter can harness entanglement to completely eliminate the uncorrelated background. Specifically, when we perform SFG with entangled biphotons, the coherent correlated peak at the original pump frequency is enhanced by the entanglement (responds linarly), whereas the incoherent background is not (responds quadratically) [35]. Thus, the ratio between the coherent signal and the incoherent noise can diverge in the limit of single entangled biphotons, leading to an enhancement of SNR that is ideally limitless. Indeed, SFG is recognized as ideal for quantum sensing applications that rely on entanglement, such as quantum radar and imaging with quantum illumination [39, 37]. Furthermore, the enhanced SNR of SFG detection directly indicates the time-energy entanglement: While the biphoton amplitude of each individual signal-idler frequency-pair in the spectrum is well below the detection threshold of $1\sqrt{\text{photons}}/\text{mode}$, the coherent sum of all modes across the spectrum can generate a well-detectable SFG signal with practically no noise background due to the ideal rejection of accidental coincidences (as we discuss hereon). Thus, efficient SFG with entangled biphotons can transform optical sensing in low-signal environments.
Unfortunately however, direct SFG at the single-photon level is ridiculously inefficient (typically $\eta_{SFG}\!=\!10^{-9}$ or lower). Consequently, SFG of entangled photons was used so far only to expose the fundamental quantum nature of time-energy entangled photons [36]), but not for practical applications of quantum sensing, where a qualitative experimental breakthrough is needed that will enhance the SFG efficiency by many orders of magnitude (ideally to $\eta_{SFG}\!=\!1$).
@@
Here we overcome this fundamental inefficiency barrier by stimulating the nonlinear interaction in the detection crystal with a strong coherent pump field, which dramatically enhances the SFG efficiency of bi-photons to near-unity [40]. Coherent stimulation—now phase-dependent—underpins nonlinear SU(1,1) interference, where the relative phase between the pump field and biphotons determines the energy flow in the second nonlinear crystal (from biphotons to pump or vice versa). Previously, this non-classical SU(1,1) interference was used for detecting biphotons (Fig. 1) by measuring the spectrum of the biphotons after the detection crystal. The resulting high-contrast interference fringes reveal the nonlinear pairwise interference of photons, with fringe contrast serving as a non-classical witness [33]. Remarkably, this process remains efficient even in lossy conditions, selectively extracting correlated biphotons. Moreover, SU(1,1) coherence enabled parametric homodyne detection [41], which uses the pump as a common phase reference to measure a specified two-mode quadrature of all signal-idler frequency-pairs across a broad spectrum. And yet, standard SU(1,1) detection does not capture the complete quantum correlation, since it still relies on local measurements of the down-converted photons.
@@
To obtain the full correlation information, we return to global detection by SFG, but we stimulate the SFG interaction coherently (see Fig. 1a). This method leverages the near unit-efficiency of SU(1,1) interference while shifting detection from the biphotons back to the pump—measuring the pump’s depletion, gain, or phase shift due to the coherent SFG conversion. In doing so, it combines SFG’s global, noise-rejecting nature with the efficiency of SU(1,1) interference.
@@
The major experimental challenge of coherent SFG is the need to measure a small SFG gain / depletion on top of an intense pump beam (of $\sim\!400$mW power), which saturates the detector by orders of magnitude. To overcome this we resort to methods of coherent detection: We mix the output beam of pump+SFG on a beam-splitter with a local oscillator at the pump frequency, whose phase is tuned to destructively interfere with the pump, thereby coherently removing most of the strong pump background from the desired SFG result prior to detection. Using coherent seeding of the SFG process and lock-in detection we are able to stabilize the phase of the small SFG signal, as explained hereon.
@@
In addition, we present an analytic quantum model (see the Methods section) of coherent SFG that we developed to fit the experimental data. The model provides a closed expression for the SFG amplitude beyond the “undepleted pump” approximation [14] by treating the three-wave mixing process as a perturbative expansion in orders of the spontaneous parametric down conversion (SPDC) amplitude. Our model is applicable in a wide range of parametric gain, from the low-gain regime of single entangled photon pairs ($g\!\ll\!1$) up to the multi-photon regime of high quadrature squeezing ($g\!\geq\!1$).
@@ 3 Experiment and Results
The main results are shown in Fig. 3, presenting the correlation in both energy-sum and time-difference (or group delay). The top panel (a) reflects the SFG spectrum of recombination, highlighting the precise match of the sum-energy of the biphotons with the original pump energy and confirming that the process is dominated by the recombination of correlated photons back to the pump frequency $\nu_{SFG}\!=\!\nu_{p}$, while the probability of accidental SFG to other SFG frequencies from non-twin photons vanishes (as theoretically predicted, see below). Specifically, panel (a) shows the RF spectrum of the of the detected pump+SFG intensity, demonstrating a single peak of high contrast (>37>\!37dB) at the lock-in dither frequency $\omega$. The measured intensity modulation is due to the beat of interference between the carrier pump at $\nu_{P}$ and the SFG sidebands at $\nu_{SFG}\!=\!\nu_{P}\!\pm\!\omega$ due to the phase modulation of the SPDC, thereby capturing the frequency-difference between the SFG signal and the pump at Hz-level resolution (2020Hz width), limited only by the integration time of the RF analyzer. The background of the SFG peak is purely the electronic noise floor of our digital spectrum analyzer (Moku:Lab). The absence of an optical contribution to the background at other frequencies $\nu\!\neq\!\omega$ indicates that the recombination of correlated biphotons dominates and that accidental SFG events are insignificant.
@@
The central panel (b) of Fig. 3 reflects the correlation in time-difference between the photons by observing the dependence of the SFG spectral peak (at the dither frequency $\omega=173$KHz) on the relative group delay between the signal and the idler photons before entering the SFG crystal. We continuously tune the group delay (and GDD) by varying the amount of material that the beam of biphotons traverses in the spectral-shaper (by shifting the Sapphire prisms in/out). The additional material dispersion imprints a quadratic spectral phase $\phi(\nu)\!=\!\beta_{2}\nu^{2}$ on the broad spectrum of spontaneous biphotons, which in time indicates a group delay (and dispersion) between each signal frequency and its twin idler.; and the SFG signal extracted from the lock-in detection is simply $I_{SFG}\!\sim\!\text{Re}[\alpha_{r}^{*}\sum_{\nu}a_{s}(\nu)a_{i}(-\nu)\mathrm{e}^{\mathrm{i}\phi(\nu)}]$, which reflects a homodyne measurement of the SFG sum over all the signal-idler pairs with $\alpha_{r}$ the coherent amplitude of the pump residue at the Sagnac output (see derivation below). Since even small GDD values alter the phase of the spontaneous SFG field, the variation of the GDD is immediately reflected in the homodyne, causing the sharp peak. The bottom panels (c) and (d) show the spectral SU(1,1) interferograms for two specific GDD values, marked A and B on the central panel (b): A - near GDD compensation (bottom right, blue), where the spectral phase is flat; and B - at a large GDD value (bottom left, violet), where the large quadratic spectral phase is imprinted on the spectral fringes.
@@
Finally, all the GDD-scans of Fig. 4 with various residue phases demonstrate a sharp peak near the zero-GDD point, which is a direct signature of the time correlation of the photon pairs. Specifically, the peak width $\delta\beta\!\approx\!4\pi/\Delta\nu^{2}\approx 30\,\mathrm{fs}^{2}$ is inversely proportional to the biphotons bandwidth of $\Delta\nu\!=\!113$ THz squared (for $\varphi_{r}\!=\!0,\pi$ the peak is split). From a physical point of view, this width is also the GDD tolerance for the photon pairs to keep coherence, which reflects their time correlation of the biphotons at $\Delta t\!=\!1/\Delta\nu\!\approx\!9$ fs. Together with the measured linewidth of 2020Hz for the SFG peak, mentioned above we obtain $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!\approx\!2\!\cdot\!10^{-13}$, showing a giant violation of the classical bound (Eq. 1). The longer-range oscillations in the GDD-scans beyond the central coherent SFG peak are mainly due to the fact that we scan the GDD, rather than the group delay, which indicates that the spectral phase at the center of the spectrum varies quadratically and not linearly across the spectrum.